How do I find the answer to 1/4(one fourth)+5/X(five over X)which equals 3/8.(three over eight)?
I dont understand this at all.
I would appreciate your help thanks.

Answer by nerdybill(7384)

1/4(one fourth)+5/X(five over X)which equals 3/8
.
1/4 + 5/x = 3/8
Multiply both sides by 8x (to get rid of denominators):
8x(1/4 + 5/x) = 8x(3/8)
2x + 40 = 3x
40 = x